OpenStudy (anonymous):

A cubical box is to be constructed from three kinds of building materials. The material used in the four sides of the box costs 2c/in^2, the material in the bottom costs 3c/in^2 and the material for the lid costs 4c/in^2. Estimate the additional total cost of all the building materials if the length of a side is increased from 20 in to 21 in. can someone help me solve this problem. Its in the marginal cost/linear approximation chapter of my calc book

OpenStudy (anonymous):

Cost: C(a) = 2 ( 4a^2) + 3a^2 + 4a^2 = 15a^2 Marginal Cost: C'(20) = 30 a a' = 30 * 20 * 1 = 600 cents

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Do you have any question?

OpenStudy (anonymous):

I actually figured it out on another thread someone kind of taught it to me. I just dont get what most of the stuff represents.